# Structure-Preserving Neural Ordinary Differential Equations for Stiff Systems

### v3

@@ 5 Numerical Examples
where the initial latent variable is projected through $\tilde{u}_{0}=f_{E}(U_{0})$. Note that the loss is computed in the original coordinates; thus, the autoencoder is optimized along with the NODE during optimization step. The inclusion of this autoencoder network is optional in our framework. Our implementation of the NODE uses Jax [23] as a backend. Automatic differentiation is used to compute gradients of the loss function with respect to the trainable parameters, and the Adam optimizer from the Optax [24] package is used for optimizations.
@@ 5.1 Taming Transient Growth
where $\sin(x)$ is interpreted to be a component-wise application of the function. The linear operator, $A$, is chosen such that in the limit of $\epsilon\to 0$ there is initial transient growth followed by long-term decay. This example is influenced by work done in [25] where the author analyzes how the norm of the matrix exponential contributes to the runtime of quantum algorithms for linear ordinary differential equations (i.e., $\epsilon=0$). Some relevant quantities used in their analysis are
@@
where $\sigma(\cdot)$ represents the the set of eigenvalues of the argument. The values of $(\alpha,\gamma)$ describe the long term behavior of the norm of $e^{At}$. We focus on the case when $\alpha<0$, which implies $\lim\limits_{t\to\infty}||e^{At}||\to 0$. While this gives the long term behavior, the quantity $\alpha$ does not give any insight into the short term behavior around $t=0$. The sign of $\gamma$ tells us whether $||e^{At}||$ has an initial grown or decay around $t=0$, $\gamma>0$ corresponding to growth and $\gamma<0$ corresponding to decay. To illustrate how $\gamma$ impacts the short timescale we consider two matrices in $\mathbb{R}^{2\times 2}$,
@@
As both matrices have eigenvalues equal to $-2$, both systems have $\alpha<0$, but $\gamma(A_{\text{decay}})<0<\gamma(A_{\text{bump}})$. This shows that $e^{tA_{\text{decay}}}$ has norm that decays for all $t>0$ and $e^{tA_{\text{bump}}}$ has an initial growth in the norm before the decay as is shown in Figure 2.
Figure 2: Plot of $||e^{tA_{\text{bump}}}||$ and $||e^{tA_{\text{decay}}}||$ versus time $t$. Blue curve corresponds to positive $\gamma$ defined in equation (22b) and the orange corresponds to negative $\gamma$.
For our experiments, we consider $A=A_{\text{bump}}$. and $\epsilon\in\{0,1.0\}$. The initial bump in the norm $||e^{tA_{\text{bump}}}||$ makes the dynamics more challenging to learn, as this implies initial transient growth before the long term decay. Data is generated using an adaptive Runge-Kutta-Fehlberg scheme. There are 1000 initial conditions evolved using 50 adaptive timesteps and a relative tolerance of $10^{-8}$. For training, only the data at a few points in time are utilized. For $\epsilon=0.0$ only the initial and final timesteps are utilized and for $\epsilon=1.0$ only timesteps $0,10,20,30,40$, and $50$ are utilized. Note that the exponential integrator can learn the dynamics for $\epsilon=0.0$ with a single timestep, but the IMEX scheme is unable to capture the dynamics. In our network, the nonlinear term is treated using a bilinear form network when $\epsilon=0.0$ and Lipschitz controlled layers when $\epsilon=1.0$. In both cases 2 layers and 100 hidden dimensions are used. The Adam optimizer from Optax is used for training with a learning rate of 0.01 and batch size of 1000. The training is run for 1000 iterations for each case. For comparison, the integration scheme is replaced with an implicit-explicit (IMEX) scheme. The predictions of the models for $\epsilon=0.0$ are shown in Figure 3 and for $\epsilon=1.0$ are shown in Figure 4. Learning with the exponential integrator yields more accurate results than the IMEX integrator for both values of $\epsilon$, the difference is more pronounced for the $\epsilon=1.0$.
@@ 5.2 Robertson’s Chemical Reaction
The training data consists of 1000 trajectories generated using MATLAB’s ODE15s solver. Each trajectory has initial condition $(1,0,0)$ plus a small perturbation. The timesteps are logarithmically spaced in the interval $[4\times 10^{-6},4\times 10^{6}]$ with 50 timesteps. The minimum and maximum timesteps are $3.03\times 10^{-6}$ and $1.72\times 10^{6}$, respectively. The model is trained for 10,000 epochs with a learing rate of 0.001 decaying by a factor of 0.99 after each epoch. In [1] it is shown that training the imbalance in the magnitude of the components can lead to gradient pathologies. The trajectory for $y_{2}$ is rescaled to $\hat{y}_{2}=10^{4}y_{2}$ so that inputs and outputs of the network are $O(1)$. To recapture the original system all that is required is scaling $\hat{y_{2}}$ by $10^{-4}$.
@@
Two NODEs are created to approximate the Robertson problem: one that utilizes the autoencoder and one that does not. All other model and training parameters remain constant. The nonlinear networks in the latent dynamics consist of 2 layers and 100 hidden dimensions. The batch size is chosen to be 100. The learning rate is initially 0.01 and reduced by a factor of 0.99 at each step. The training is run for 10,000 epochs in each model. Lipschitz controlled layers are used to represent the nonlinear operator. In Figure 5, predictions for both models with initial condition $(1,0,0)$ are displayed. It is clear from this figure that the application of the autoencoder produces an accurate model for all components. Here the autoencoder acts like a coordinate transformation. We stress that both versions of the model are able to robustly integrate using timesteps ranging from $3.03\times 10^{-6}$ to $1.72\times 10^{6}$ without rescaling time. In both cases, the model is learning to extract a stiff linear operator that enables efficient integration on all scales.
@@
Figure 5: Top: Results of training without an autoencoder. Bottom: Results of training with an autoencoder. Dotted lines represent the true solution and colored solid lines represent the prediction given by the learned model.
@@
Many existing approaches for learning stiff dynamics such as the Robertson problem rely on reparameterizing time and transforming the system into a non-autonomous system. In [14] the authors resolve the stiffness by learning a reparameterization in time. The resulting learned dynamical system is a nonstiff surrogate and may be cheaply evaluated using an explicit integrator. In data driven approaches for stiff dynamics [13, 26] a fixed time transformation is used, such as $\Delta\tilde{t}=-\frac{1}{\log{\Delta t}}$. Our proposed approach avoids the need to know the time scale change a prior or learning the time reparameterization. The proposed approach can therefore bridge large variations in time scales–a desirable property for multi-scale problems that are often found in plasma physics and fluid dynamics. The only other approach that we are aware of learning directly in the original time domain is [27], in which a multiscale structure is explicitly built in neural ODE.
@@ 5.3 Kuramoto-Sivashinsky Equation
We consider the dataset used in [2]. Solutions are found by performing a Galerkin projection onto Fourier modes and using exponential time differencing to evolve the ODE forward in time [28]. After solving in Fourier space, we then transform the data back to physical space for training the neural ODEs. We sample the data at every 1.0 time units. The training data is constructed by separating the long trajectory into trajectories with 8 steps. Our batch size is 2000 and our learning rate is 0.001. The learning rate is reduced by a factor of 0.99 at each step. For this problem we use a Lipschitz controlled feed-forward neural network for the nonlinear part using 2 layers and inner dimension 200. We observe that Lipschitz control simultaneously improves training performance and the long-term stability of the dynamic model.
There are a few additional steps taken to train this model effectively. First, a shift of the spectrum to the Hurwitz matrix is applied during training. This is required since the KS system has both positive and negative eigenvalues and the Hurwitz parametrization requires the real part of the eigenvalues to be negative. Second, an autoencoder that projects the dynamics down to a latent space. There are many attractive options for model reduction. A popular method is utilizing projection based reduced-order modeling, which attempt to find a low-dimensional trial subspace that can represent the state space. Most of these trial subspaces are linear and can not capture the complicated dynamics of the KS equation. Moreover, the KS equation has a slow decaying Kolomogorov $n$-width [29], which implies linear trial subspaces may not accurately represent the system. The power behind autoencoders is that they attempt to find an identity mapping with fewer dimensions. If the data can be represented in a lower dimensional nonlinear manifold, then optimization of the encoder and decoder finds the optimal low-order representation. For this example 64 grid points are projected into a latent space of dimension 20 using an autoencoder. For $L=22$, the KS equation has 3 positive Lyapunov exponents, resulting in the Kaplan-York dimension of its attractor to be 5.198 [30]. Therefore, it needs an inertial manifold of at least 6 dimensions. Our second model utilizes a latent space of 7 dimensions. It is unclear if it is possible to successfully train with a lower dimensional latent space. We note that relative to our baseline architecture without an autoencoder, we observe factors of 12 and 24 speedup for the models with dimension 20 and 7, respectively.
Figure 6: Displayed is the true solution ($u_{true}$), the model prediction with latent dimension 20 and 7 ($u_{20}$,$u_{7}$) and the errors ($\rm{error}_{20},\rm{error}_{7})$

### v4

@@ 5 Numerical Examples
where the initial latent variable is projected through $\tilde{u}_{0}=f_{E}(U_{0})$. Note that the loss is computed in the original coordinates; thus, the autoencoder is optimized along with the NODE during optimization step. The inclusion of this autoencoder network is optional in our framework. Our implementation of the NODE uses Jax jax as a backend. Automatic differentiation is used to compute gradients of the loss function with respect to the trainable parameters, and the Adam optimizer from the Optax optax package is used for optimizations.
@@ 5.1 Taming Transient Growth
where $\sin(x)$ is interpreted to be a component-wise application of the function. The linear operator, $A$, is chosen such that in the limit of $\epsilon\to 0$ there is initial transient growth followed by long-term decay. This example is influenced by work done in krovi2023improved where the author analyzes how the norm of the matrix exponential contributes to the runtime of quantum algorithms for linear ordinary differential equations (i.e., $\epsilon=0$). Some relevant quantities used in their analysis are
@@
[…]
@@
As both matrices have eigenvalues equal to 2-2, both systems have $\alpha<0$, but $\gamma(A_{\text{decay}})<0<\gamma(A_{\text{bump}})$. This shows that $e^{tA_{\text{decay}}}$ has norm that decays for all $t>0$ and $e^{tA_{\text{bump}}}$ has an initial growth in the norm before the decay as is shown in Figure 2.
Figure 2: Plot of $e^{tA_{\text{bump}}}||$ and $e^{tA_{\text{decay}}}||$ versus time $t$. The blue curve corresponds to positive $\gamma$ defined in equation (22b) and the orange corresponds to negative $\gamma$.
For our experiments, we consider $A=A_{\text{bump}}$. and $\epsilon\in\{0,1.0\}$. The initial bump in the norm $e^{tA_{\text{bump}}}||$ makes the dynamics more challenging to learn, as this implies initial transient growth before the long term decay. Data is generated using an adaptive Runge-Kutta-Fehlberg scheme. There are 1000 initial conditions evolved using 50 adaptive timesteps and a relative tolerance of $10^{-8}$. For training, only the data at a few points in time are utilized. For $\epsilon=0.0$ only the initial and final timesteps are utilized and for $\epsilon=1.0$ only timesteps $0,10,20,30,40$, and $50$ are utilized. Note that the exponential integrator can learn the dynamics for $\epsilon=0.0$ with a single timestep, but the IMEX scheme is unable to capture the dynamics. In our network, the nonlinear term is treated using a bilinear form network when $\epsilon=0.0$ and Lipschitz controlled layers when $\epsilon=1.0$. In both cases 2 layers and 100 hidden dimensions are used. The Adam optimizer from Optax is used for training with a learning rate of 0.01 and batch size of 1000. The training is run for 1000 iterations for each case. For comparison, the integration scheme is replaced with the implicit-explicit (IMEX) SSP2(2,2,2) L-Stable Scheme defined in Table II of pareschi. The predictions of the models for $\epsilon=0.0$ are shown in Figure 3 and for $\epsilon=1.0$ are shown in Figure 4. Learning with the exponential integrator yields more accurate results than the IMEX integrator for both values of $\epsilon$, the difference is more pronounced for the $\epsilon=1.0$.
@@ 5.2 Robertson’s Chemical Reaction
The training data consists of 1000 trajectories generated using MATLAB’s ODE15s solver. Each trajectory has initial condition $(1,0,0)$ plus a small perturbation. The timesteps are logarithmically spaced in the interval $4\times 10^{-6},4\times 10^{6}]$ with 50 timesteps. The minimum and maximum timesteps are $3.03\times 10^{-6}$ and $1.72\times 10^{6}$, respectively. The model is trained for 10,000 epochs with a learing rate of 0.001 decaying by a factor of 0.99 after each epoch. In kim2021stiff it is shown that training the imbalance in the magnitude of the components can lead to gradient pathologies. The trajectory for $y_{2}$ is rescaled to $\hat{y}_{2}=10^{4}y_{2}$ so that inputs and outputs of the network are $O(1)$. To recapture the original system all that is required is scaling $\hat{y_{2}}$ by $10^{-4}$.
@@
Two NODEs are created to approximate the Robertson problem: one that utilizes the autoencoder and one that does not. All other model and training parameters remain constant. The nonlinear networks in the latent dynamics consist of 2 layers and 100 hidden dimensions. The batch size is chosen to be 100. The learning rate is initially 0.01 and reduced by a factor of 0.99 at each step. The training is run for 10,000 epochs in each model. Lipschitz controlled layers are used to represent the nonlinear operator. In Figure 5(a), predictions for both models with initial condition $(1,0,0)$ are displayed. It is clear from this figure that the application of the autoencoder produces an accurate model for all components. Here the autoencoder acts like a coordinate transformation. The latent space dynamics are presented in Figure 5(b), illustrating how the non-trivial transformation induced by the autoencoder shifts the stiffness of the original system from the nonlinear operator into the linear operator. A non-trivial, stiff, latent dynamics is learned under a large temporal variation. We stress that both versions of the model are able to robustly integrate using timesteps ranging from $3.03\times 10^{-6}$ to $1.72\times 10^{6}$ without rescaling time. In both cases, the model is learning to extract a stiff linear operator that enables efficient integration on all scales.
@@
(a) Top: training without a transformation; Bottom: with an transformation. Dotted lines are the true solution; colored solid lines are model predictions.
(b) Latent space dynamics for Robertson’s chemical reaction.
Figure 5: (a) Prediction vs. truth with and without a transformation. (b) Latent dynamics corresponding to the same system when transformation is utilized.
@@
Many existing approaches for learning stiff dynamics such as the Robertson problem rely on reparameterizing time and transforming the system into a non-autonomous system. In caldana2024neural the authors resolve the stiffness by learning a reparameterization in time. The resulting learned dynamical system is a nonstiff surrogate and may be cheaply evaluated using an explicit integrator. In data driven approaches for stiff dynamics xie2024latent, nockolds2025constant a fixed time transformation is used, such as $\Delta\tilde{t}=-\frac{1}{\log{\Delta t}}$. Our proposed approach avoids the need to know the time scale change a prior or learning the time reparameterization. The proposed approach can therefore bridge large variations in time scales–a desirable property for multi-scale problems that are often found in plasma physics and fluid dynamics. We observe that our approach performs better than caldana2024neural with a much smaller error for the same test, particularly in the initial stage of the dynamics. caldana2024neural obtains a worse result for this particular case due to its working on a rescaled dynamics, which does not prioritize the error in the true dynamics. The only other approach that we are aware of learning directly in the original time domain is chen2025due, in which a multiscale structure is explicitly built in neural ODE.
@@
Figure 6: Model and true trajectories with Gaussian noise scaled by 0.01 (top) and 0.1 (bottom).
5.2.1 Network Ablation Studies
To evaluate the robustness of the proposed network, we perform two ablation studies: (1) assessing the model’s sensitivity to noisy inputs and (2) analyzing the effect of the Lipschitz constraint on the nonlinear operator. In the first study, Gaussian noise is added to each trajectory used for training, with noise levels scaled by factors of 0.01 and 0.1. As shown in Figure 6, the model remains robust across all noise amplitudes, demonstrating strong resilience to perturbations in the training data. Table 1 presents the results of the second study, which investigates the impact of varying the Lipschitz constant on model performance. Three configurations are compared: no constraint, $L=1.0$, and $L=2.0$. Additionally, the scaling factor applied to the initial weights of the nonlinear function is varied from $10^{-3}$ to $10^{0}$. The results indicate that imposing a Lipschitz constraint significantly improves accuracy and stability across all initialization scales. In contrast, models trained without the constraint exhibit less consistent and occasionally unstable behavior. These findings also highlight that the performance remains sensitive to the magnitude of the initial weights, underscoring the importance of proper initialization when applying Lipschitz regularization. In each case the learning rate is initialized at 0.01 and has a decay rate of 0.9. The network has 100 hidden dimensions and 2 layers. A total of 20000 iterations are done and the final loss is reported in Table 1.
Table 1: Effect of Lipschitz control constant and initilzation scale on model performance. Values represent the final training loss across different initialization scales.
Init. Scale No Lipschitz $L=1.0$ $L=2.0$
$10^{-3}$ 0.052662 0.004060 0.004084
$10^{-2}$ 0.413053 0.004100 0.004087
$10^{-1}$ 0.004639 0.000201 0.000261
$10^{0}$ 0.023333 0.004182 0.004027
@@ 5.3 Kuramoto-Sivashinsky Equation
We consider the dataset used in linot2023stabilized. Solutions are found by performing a Galerkin projection onto Fourier modes and using exponential time differencing to evolve the ODE forward in time kassam2005fourth. After solving in Fourier space, we then transform the data back to physical space for training the neural ODEs. We sample the data at every 1.0 time units. The training data is constructed by separating the long trajectory into trajectories with 8 steps. Our batch size is 2000 and our learning rate is 0.001. The learning rate is reduced by a factor of 0.99 at each step. For this problem we use a Lipschitz controlled feed-forward neural network for the nonlinear part using 2 layers and inner dimension 200. We observe that Lipschitz control simultaneously improves training performance and the long-term stability of the dynamic model.
There are a few additional steps taken to train this model effectively. First, a shift of the spectrum to the Hurwitz matrix is applied during training. This is required since the KS system has both positive and negative eigenvalues and the Hurwitz parametrization requires the real part of the eigenvalues to be negative. Second, an autoencoder that projects the dynamics down to a latent space. There are many attractive options for model reduction. A popular method is utilizing projection based reduced-order modeling, which attempt to find a low-dimensional trial subspace that can represent the state space. Most of these trial subspaces are linear and can not capture the complicated dynamics of the KS equation. Moreover, the KS equation has a slow decaying Kolomogorov $n$-width mojgani2023kolmogorov, which implies linear trial subspaces may not accurately represent the system. The power behind autoencoders is that they attempt to find an identity mapping with fewer dimensions. If the data can be represented in a lower dimensional nonlinear manifold, then optimization of the encoder and decoder finds the optimal low-order representation. For this example 64 grid points are projected into a latent space of dimension 20 using an autoencoder. For $L=22$, the KS equation has 3 positive Lyapunov exponents, resulting in the Kaplan-York dimension of its attractor to be 5.198 edson2019lyapunov. Therefore, it needs an inertial manifold of at least 6 dimensions. Our second model utilizes a latent space of 7 dimensions. This finding is consistent with the previous work linot2020deep, where the authors found that for the exact same KS test, the minimal latent dimension for their model to train reasonably is 7. We note that relative to our baseline architecture without an autoencoder, we observe factors of 12 and 24 speedup for the models with dimension 20 and 7, respectively.
Figure 7: Displayed is the true solution ($u_{true}$), the model prediction with latent dimension 20 and 7 ($u_{20}$$u_{7}$) and the errors ($\rm{error}_{20},\rm{error}_{7})$ .
@@
6 Discussions
A critical component of solving high-dimensional problems is dimensionality reduction. We pursue this through an autoencoder that compresses the state variables into a latent representation whose evolution satisfies an ordinary differential equation. Our approach can be viewed as an extension of latent neural ODEs rubanova2019latent, but with additional structure-preserving mechanisms built into the latent dynamics. Latent neural ODEs learn a low-dimensional latent space in which continuous-time dynamics evolve, enabling the modeling of irregularly sampled or partially observed trajectories. Our framework extends this idea by introducing Hurwitz-stable linear operators, Lipschitz-controlled nonlinear components, and exponential integration, ensuring stable and efficient learning even for stiff or multiscale systems. This structure makes the method particularly suitable for physical systems where long-term accuracy and numerical robustness are essential. Beyond latent ODEs, there are other extensions of the neural ODE framework designed to enhance representational capacity. Augmented neural ODEs dupont2019 expand the state space with additional coordinates, allowing the model to represent more complex trajectories and overcome topological limitations of standard NODEs. In our examples, which involve first-order systems solved directly for the full state variable, augmentation is not physically motivated. However, the principles of exponential integration, Hurwitz stability, and Lipschitz control can naturally extend to augmented formulations. For PDEs in multiple spatial dimensions, the number of degrees of freedom in the discretization grows rapidly with the spatial dimension. To address such large-scale problems, it may be necessary to employ sparse or localized operators, such as convolutional, patch-based, or graph-based embeddings, to perform efficient coarse-graining before applying the learned structure-preserving neural ODE. Extending these ideas to high-dimensional PDEs represents an important step toward modeling realistic multi-scale physical phenomena, including fluid turbulence, plasma dynamics, and climate processes.
The proposed framework shows potential for extension to more complex dynamical systems beyond the examples considered in this work. In particular, it can naturally be applied to multi-scale systems, where components evolve on widely separated time scales. In our recent work on fast-slow dynamical systems serino2024intelligent, we demonstrated that a structure-preserving formulation can accurately capture the fast, dissipative dynamics through a linear-nonlinear split while maintaining stability and efficiency. The present approach builds on these ideas, and can be directly applied to the fast subsystem to achieve stable integration across multiple scales, while future work will target systems that exhibit strongly oscillatory or resonant behavior. For non-smooth systems, such as those involving shocks, contact discontinuities, or material interfaces, the performance of the method will depend primarily on the spatial discretization employed. Nevertheless, the exponential time-stepper introduced in this paper can be readily combined with shock-capturing or discontinuity-aware spatial schemes, enabling robust integration even in the presence of non-smooth features. We also note that hybrid dynamical systems, which involve discrete-continuous interactions or state-dependent switching, fall outside the continuous ODE framework considered here. Extending the current formulation to handle event-driven updates or discontinuous vector fields would require additional mechanisms, such as switching-aware integration or piecewise structure-preserving embeddings. Addressing these challenges represents a compelling direction for future research toward applying structure-preserving neural ODEs to a broader class of real-world, multi-physics problems.
@@ 7 Conclusions
While our experiments focus on canonical stiff systems, the proposed framework is general and can be applied to other domains. In particular, neural ODEs trained on image, speech, or sequential data may exhibit effective stiffness during optimization, making our structure-preserving approach potentially beneficial for broader deep learning tasks. We plan to investigate this connection in future work by evaluating the method on standard benchmarks where stiffness may arise implicitly during training. Beyond deep learning, the framework can also be extended to real-world scientific and engineering problems such as collisional-radiative modeling, climate and geophysical forecasting, and computational fluid dynamics, where stiffness and multi-scale interactions are prevalent. These applications represent natural testbeds for assessing the scalability, stability, and interpretability advantages of structure-preserving neural ODEs in realistic, high-dimensional environments.
@@ Acknowledgments
Data Availability
The code and datasets analyzed during the current study are available from the corresponding author on reasonable request.